Displacement y of particle is given by y=2t+t2-2t2.the velocity of the particle when acceleration is zero

Grade:12

2 Answers

Vikas TU
7 years ago
velocity of the ;particle is the differential of the displacent of y.
.i.e. dy/dt = 2 + 2t – 6t^2
acceleration is the differentiation of velocity or double differentale of displacement.
.i.e. a = 2 – 12t = 0 
t = 1/6
velocity at t = 1/6 now u can calcclate by putting t in velocity eqn.

Note: that last term is 2t^3 i think not t^2. 
Anuj Rampuriya
2 years ago
velocity of the particle is the differential of y
 
now,
          dy/dt=2t-6t^2
          dy/dt={2-12t}-eq 1
           
          0=2-12t
          12t=2
           t=2/12
           t=1/6
putting value of t in the given eq 
when acceleration is zero  
v=2+2/6-6(1/6)2
=2+1/3-1/6
=13/6